An aeroplane takes off at an angle of 60o to the horizontal. If the component of the velocity along the horizontal is 200m/s, what is the actual velocity? Also find the vertical component of its velocity
1
Expert's answer
2018-08-15T09:00:08-0400
The component of the velocity along the horizontal is
v_x=v cos⁡α.
The actual velocity is
v=v_x/cos⁡α =200/cos⁡60 =200/0.5=400 m/s.
The vertical component of its velocity:
v_y=v sin⁡α=400 sin⁡60=346 m/s.